[figure: 9.7 cm]

ELECTRONIC STRUCTURE AND MAGNETIC PROPERTIES

OF THE GdRhAl TERNARY COMPOUND

J. Jarosz, E. Talik Institute of Physics, University of Silesia, ul. Uniwersytecka 4, 40- 007 Katowice, Poland

A new ternary compound GdRhAl was obtained using the Czochralski method from a levitated melt. The temperature dependence of the electrical resistivity, magnetic susceptibility, electron spin resonance and the XPS investigation were performed. The influence of the Rh 4d band contribution to the density of states at the Fermi level on the magnetic and transport properties of the compound was observed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : J.Jarosz

Address for correspondence : ul. Uniwersytecka 4 40-007 Katowice

Email address : jjarosz@us.edu.pl